# A Control Allocation Algorithm for Hypersonic Glide Vehicles with Input Limitations

### v2

@@ Abstract
$B_{\nu_{x}},B_{\nu_{y}},B_{\nu_{z}}$  = Conditionalized rows of $\mathbf{B}$ associated with roll, pitch, and yaw
@@ 1 Introduction
In recent years, hypersonic glide vehicles (HGVs) have received increasing attention in academic and industrial research and development [1]. These vehicles, designed to navigate through lower atmospheric layers at hypersonic speeds, present unique challenges that require precise control for optimal performance and guaranteed stability. Some HGV designs incorporate redundant control effectors, necessitating a control allocation system capable of distributing desired forces and moments among a redundant set of control effectors to achieve the desired vehicle response [2].
Traditional control allocation methods for overactuated aerial vehicles have evolved significantly. For most aerospace applications, the pseudoinverse-based control allocation (PICA) method [3], utilizing the Moore-Penrose inverse, is commonly used due to its simplicity. However, PICA does not adequately account for control input limitations. An alternative and more precise approach is to formulate the constrained control allocation problem as a quadratic programming (QPCA) problem and solve it using a QP solver [4, 5]. This method can accurately handle asymmetric input constraints and coupling effects between control effectors. However, the computational intensity of QP solvers poses significant challenges for real-time implementation on systems with limited computational resources, such as embedded systems in hypersonic reentry vehicles [6]. The high computational demand makes it difficult to guarantee real-time performance, which is essential for the fast dynamics of HGVs. To address the need for a computationally less complex solution, Bordignon introduced the redistributed pseudoinverse control allocation (RPICA) method [7]. RPICA iteratively approximates feasible solutions by redistributing residuals over the attainable moment set (AMS), offering a balance between computational efficiency and practicality. A scaled variant, RSPICA, was later proposed to maintain the directionality of the input vector in multiple-input, multiple-output (MIMO) systems under saturation conditions. Despite advancements in computational simplicity, both RPICA and RSPICA face challenges when dealing with systems characterized by asymmetric control input limits and strong inter-directional coupling between control effectors. This issue is particularly evident in certain classes of HGVs where only positive control deflections are permissible. The inability of these methods to effectively handle asymmetric constraints and coupling effects limits their applicability in such scenarios.
@@
Over the past few years, the German Aerospace Center (DLR) has undertaken a substantial research initiative focused on the development of the Generic Hypersonic Glide Vehicle 2 (GHGV-2). This research aims to comprehensively explore the capabilities and performance characteristics of strategic hypersonic glide vehicles, shedding light on their implications for understanding hypersonic threats [8, 9]. The GHGV-2 adopts a waverider design to optimize lift-to-drag ratios at high Mach numbers [10]. The GHGV-2 is depicted in Figure 1, with an external view shown in Figure 1(a) and a sectional view highlighting relevant subsystems in Figure 1(b). Key subsystems include the thermal protection system (TPS), guidance, navigation, and control (GNC) system, battery, and actuators. During endoatmospheric flight, the GHGV-2 utilizes four integrated flaps as control effectors to regulate its attitude. Small propulsors are integrated in exoatmospheric conditions to generate the necessary moments for attitude adjustments.
The GHGV-2 is launched via a multi-stage rocket booster as part of the investigated mission profile. After ignition and acceleration, the vehicle separates from the launch system at an altitude of approximately 100 km, initiating a parabolic re-entry trajectory. Once sufficient dynamic pressure is achieved during descent, the vehicle transitions to aerodynamic control using four surfaces,two located on the upper and two on the lower side. Following re-entry, the GHGV-2 tracks a predefined flight path angle to sustain its maximum lift-to-drag ratio en route to the target location. Due to sustained hypersonic speeds, the vehicle is exposed to significant thermal loads during this atmospheric glide phase. Consequently, any viable control allocation algorithm must generate the necessary control moments for maneuvering and account for thermal effects, ideally minimizing heat loads to reduce the burden on the TPS and lower the infrared signature, a key consideration for military applications. Another critical challenge arises from state-dependent control input limitations stemming from the wide range of operating conditions. Increased dynamic pressure at lower altitudes and higher speeds restricts actuator authority compared to flight conditions involving lower speeds or air densities. This variability demands a control allocation approach explicitly incorporating state-dependent constraints to ensure robust actuator performance throughout the mission.
@@ 2 The General Control Allocation Problem
where $\mathbf{x}(t)\in\mathbb{R}^{n}$, $\mathbf{u}(t)\in\mathbb{R}^{m}$ and $\mathbf{y}(t)\in\mathbb{R}^{p}$. The function $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ represents a Lipschitz continuous system function, $\mathbf{h}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ a Lipschitz continuous output function, and $\mathbf{g}(\mathbf{x}(t))\in\mathbb{R}^{n\times m}$ is referred to as the control effectiveness matrix. In the field of flight control, inversion-based control methodologies, such as NDI, are commonly used to impose desired closed-loop behavior on the MIMO dynamics [11]. To do so, the time derivatives of the introduced nonlinear dynamics from Eq. (1) and Eq. (2) are derived and reformulated using the Lie derivative notation [12]:
@@
A relationship between the input and the output can be established if $\mathbf{G}(\mathbf{x}(t))\neq 0$. For cases in which $\mathbf{G}(\mathbf{x}(t))=0$, the relationship must be further derived until the system input affects the system output. The number of time derivations needed is called the relative degree and is represented by $r$ [13]. For the sake of brevity, we consider here an example of a first-order system. For such a system with a relative degree of one, the NDI-based control law can be stated as:
@@
with $\bm{\nu}(t)\in\mathbb{R}^{o}$ being a virtual control input vector created by a linear controller to impose the desired closed-loop characteristics on the controlled variables such that $\bm{\nu}(t)=-\mathbf{K}\mathbf{y}(t)$, with $\mathbf{K}$ being a Hurwitz gain matrix. It is evident that the inversion of the control effectiveness matrix $\mathbf{G}(\mathbf{x}(t))$, from now on for simplicity referred to as $\mathbf{B}$, maps a desired $\bm{\nu}(t)$ into the control space $\mathbb{R}^{m}$. In cases in which $m=n$ and $\operatorname{rank}(\mathbf{B})=n$, this leads to a unique solution. In the case of over-actuation, where $m>n$, the inversion of the matrix is not easily available, since the mapping into the control space is not unique and, mathematically speaking, under-determined. In the flight control literature this problem is often referred to as the control allocation problem [14]. In general, the control allocation problem can be interpreted as a constrained optimization problem since magnitude and rate limitations on each actual available control effector $u_{i}(t)$ need to be considered. The rate limits are often defined as symmetric lower and upper bounds for each actuator, leading to the following rate saturation function for each actuator $i$:
@@
The described admissible control input space for a static toy example with two magnitude- and rate-constrained control inputs $u_{1}$ and $u_{2}$ is given in Fig. 2.
@@
In the context of flight control, the resulting convex set $\bm{D}$ is often referred to as the attainable moment set (AMS) [15]. An example of an AMS, in which dynamic effects coming from the rate limits are neglected, for a hypersonic glide vehicle with four control effectors ($m=4$) is shown in Fig. 3.
@@
with $p$ being a real number used to define the considered norm. In this work and without further discussion, the $\ell_{2}$-norm was chosen. Approaches based on the $\ell_{1}$-norm or the $\ell_{\infty}$-norm could also lead to desirable results; interested readers are referred to [16, 17, 18, 19, 14] for deeper discussions on the topic of the influence of different norms on the control allocation performance. Based on the decision to minimize the $\ell_{2}$-norm of the least-square problem, the optimization problem can be formulated as a quadratic programming control allocation (QPCA) problem [4, 5].
2.1 Control Allocation Algorithms
@@ 2.1.1 Pseudoinverse-Based Control Allocation
The majority of the control allocation problems are solved by pseudoinverse-based control allocation (PICA) using the Moore-Penrose pseudoinverse, which provides the min-norm solution and which implicitly solves the following unconstrained least-square problem for the allocation of the incremental virtual control input vector $\bm{\nu}$:

### v3

@@ Abstract
$\mathbf{B}_{\nu_{x}},\mathbf{B}_{\nu_{y}},\mathbf{B}_{\nu_{z}}$  = Conditionalized rows of $\mathbf{B}$ associated with roll, pitch, and yaw
@@ 1 Introduction
In recent years, hypersonic glide vehicles (HGVs) have received increasing attention in academic and industrial research and development [Slayer_2022]. These vehicles, designed to navigate through lower atmospheric layers at hypersonic speeds, present unique challenges that require precise control for optimal performance and guaranteed stability. Some HGV designs incorporate redundant control effectors, necessitating a control allocation system capable of distributing desired forces and moments among a redundant set of control effectors to achieve the desired vehicle response [Autenrieb_2023].
Traditional control allocation methods for overactuated aerial vehicles have evolved significantly [Oppenheimer_2006]. For most aerospace applications, the pseudoinverse-based control allocation (PICA) method [Bordignon_2002], utilizing the Moore-Penrose inverse, is commonly used due to its simplicity. However, PICA does not adequately account for control input limitations. An alternative and more precise approach is to formulate the constrained control allocation problem as a quadratic programming (QPCA) problem and solve it using a QP solver [Johansen_2004, Ola2004]. This method can accurately handle asymmetric input constraints and coupling effects between control effectors. However, the computational intensity of QP solvers poses significant challenges for real-time implementation on systems with limited computational resources, such as embedded systems in hypersonic reentry vehicles [Burken_2001]. The high computational demand makes it difficult to guarantee real-time performance, which is essential for the fast dynamics of HGVs. To address the need for a computationally less complex solution, Bordignon introduced the redistributed pseudoinverse control allocation (RPICA) method [Bordignon_1996]. RPICA iteratively approximates feasible solutions by redistributing residuals over the attainable moment set (AMS), offering a balance between computational efficiency and practicality. A scaled variant, redistributed scaled pseudoinverse control allocation (RSPICA), was later proposed to maintain the directionality of the input vector in multiple-input, multiple-output (MIMO) systems under saturation conditions [Hafner_2025]. Despite advancements in computational simplicity, both RPICA and RSPICA face challenges when dealing with systems characterized by asymmetric control input limits and strong inter-directional coupling between control effectors. This issue is particularly evident in certain classes of HGVs where only positive control deflections are permissible. The inability of these methods to effectively handle asymmetric constraints and coupling effects limits their applicability in such scenarios.
@@
Over the past few years, the German Aerospace Center (DLR) has undertaken a substantial research initiative focused on the development of the Generic Hypersonic Glide Vehicle 2 (GHGV-2). This research aims to comprehensively explore the capabilities and performance characteristics of strategic hypersonic glide vehicles, shedding light on their implications for understanding hypersonic threats [Autenrieb2021, Autenrieb2022]. The GHGV-2 adopts a waverider design to optimize lift-to-drag ratios at high Mach numbers [Gruhn2020]. The GHGV-2 is depicted in Figure 1, with an external view shown in Figure 1(a) and a sectional view highlighting relevant subsystems in Figure 1(b). Key subsystems include the thermal protection system (TPS), guidance, navigation, and control (GNC) system, battery, and actuators. During endoatmospheric flight, the GHGV-2 utilizes four integrated flaps as control effectors to regulate its attitude. Small propulsors are integrated in exoatmospheric conditions to generate the necessary moments for attitude adjustments.
The GHGV-2 is launched via a multi-stage rocket booster as part of the investigated mission profile. After ignition and acceleration, the vehicle separates from the launch system at an altitude of approximately 100 km, initiating a parabolic re-entry trajectory. Once sufficient dynamic pressure is achieved during descent, the vehicle transitions to aerodynamic control using four surfaces,two located on the upper and two on the lower side. Following re-entry, the GHGV-2 tracks a predefined flight path angle to sustain its maximum lift-to-drag ratio en route to the target location. Due to sustained hypersonic speeds, the vehicle is exposed to significant thermal loads during this atmospheric glide phase. Consequently, any viable control allocation algorithm must generate the necessary control moments for maneuvering and account for thermal effects, ideally minimizing heat loads to reduce the burden on the TPS and lower the infrared signature, a key consideration for military applications. Another critical challenge arises from state-dependent control input limitations stemming from the wide range of operating conditions. Increased dynamic pressure at lower altitudes and higher speeds restricts actuator limits compared to flight conditions involving lower speeds or air densities. This variability demands a control allocation approach explicitly incorporating state-dependent constraints to ensure robust actuator performance throughout the mission.
@@ 2 The General Control Allocation Problem
where $\mathbf{x}(t)\in\mathbb{R}^{n}$, $\mathbf{u}(t)\in\mathbb{R}^{m}$ and $\mathbf{y}(t)\in\mathbb{R}^{p}$. The function $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ represents a Lipschitz continuous system function, $\mathbf{h}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ a Lipschitz continuous output function, and $\mathbf{g}(\mathbf{x}(t))\in\mathbb{R}^{n\times m}$ is referred to as the control effectiveness matrix. In the field of flight control, inversion-based control methodologies, such as NDI, are commonly used to impose desired closed-loop behavior on the MIMO dynamics [Holzapfel_2004]. To do so, the time derivatives of the introduced nonlinear dynamics from Eq. (1) and Eq. (2) are derived and reformulated using the Lie derivative notation [Lombaerts_2012]:
@@
A relationship between the input and the output can be established if $\mathbf{G}(\mathbf{x}(t))\neq 0$. For cases in which $\mathbf{G}(\mathbf{x}(t))=0$, the relationship must be further derived until the system input affects the system output. The number of time derivations needed is called the relative degree and is represented by $r$ [Bhardwaj_2021]. For the sake of brevity, we consider here an example of a first-order system. For such a system with a relative degree of one, the NDI-based control law can be stated as:
@@
with $\bm{\nu}(t)\in\mathbb{R}^{o}$ being a virtual control input vector created by a linear controller to impose the desired closed-loop characteristics on the controlled variables such that $\bm{\nu}(t)=-\mathbf{K}\mathbf{y}(t)$, with $\mathbf{K}$ being a Hurwitz gain matrix. It is evident that the inversion of the control effectiveness matrix $\mathbf{G}(\mathbf{x}(t))$, from now on for simplicity referred to as $\mathbf{B}$, maps a desired $\bm{\nu}(t)$ into the control space $\mathbb{R}^{m}$. In cases in which $m=n$ and $\operatorname{rank}(\mathbf{B})=n$, this leads to a unique solution. In the case of over-actuation, where $m>n$, the inversion of the matrix is not easily available, since the mapping into the control space is not unique and, mathematically speaking, under-determined. In the flight control literature this problem is often referred to as the control allocation problem [Petersen_2005]. In general, the control allocation problem can be interpreted as a constrained optimization problem since magnitude and rate limitations on each actual available control effector $u_{i}(t)$ need to be considered. The rate limits are often defined as symmetric lower and upper bounds for each actuator, leading to the following rate saturation function for each actuator $i$:
@@
The described admissible control input space for a static toy example with two magnitude- and rate-constrained control inputs $u_{1}$ and $u_{2}$ is given in Figure 2.
@@
In the context of flight control, the resulting convex set $\bm{D}$ is often referred to as the attainable moment set (AMS) [Pfeile2021]. An example of an AMS, in which dynamic effects coming from the rate limits are neglected, for a hypersonic glide vehicle with four control effectors ($m=4$) is shown in Figure 3.
@@
with $p$ being a real number used to define the considered norm. The choice of the norm has a direct influence on the resulting control effort distribution and how the available actuator suite is used. In later sections of this work, the $\ell_{2}$-norm is adopted to promote a balanced distribution of control effort for the GHGV-2 and to enable an efficient least-squares formulation of the control allocation problem. Alternative formulations based on the $\ell_{1}$-norm or the $\ell_{\infty}$-norm are in many application cases equally well suited to handle actuator limitations, coupling effects, and additional preference criteria, and have been successfully applied and tested in similar flight control applications [Schierman_2004]. Nevertheless, in order to provide a more comprehensive context for the choice of the $\ell_{2}$-norm adopted in this work, the influence of different norm selections on the solution of the control allocation problem is discussed in the following subsection.
2.1 The role of norms in control allocation
By interpreting the optimization problem from a geometric perspective, different norms induce different level sets in the control input space, which interact with the admissible control input set $\bm{U}$ and the AMS $\bm{D}$ in distinct ways. As a result, the choice of norm implicitly encodes a preference for certain control distribution strategies and affects how a demanded virtual control input $\bm{\nu}(t)$ is distributed among the available actuators.
The minimization of the Euclidean $\ell_{2}$-norm corresponds to a least-squares formulation and leads to a quadratic programming (QP) control allocation problem. The $\ell_{2}$-norm promotes a smooth and balanced distribution of control effort across all available actuators and yields a unique solution whenever the feasible set is non-empty [harkegaard2003]. Due to its favorable numerical properties and balanced way of distributing the control effort within the available actuators, $\ell_{2}$-based control allocation has become a standard choice in many flight control applications [Johansen_2004, Ola2004]. In contrast, the $\ell_{1}$-norm promotes sparsity in the control input vector and tends to concentrate control effort on a reduced subset of actuators. This property can be advantageous in systems where certain effectors are preferred due to efficiency or authority. Control allocation problems based on the $\ell_{1}$-norm can be formulated as linear programs and efficiently solved using simplex-based or interior-point methods. Linear programming (LP) control allocation approaches have been successfully demonstrated in flight-tested applications, including re-entry vehicles with actuator constraints [Schierman_2004, Ikeda_2000, Bolender_2004]. The $\ell_{\infty}$-norm minimizes the maximum absolute actuator deflection within the actuator suite and can be interpreted as a worst-case input allocation strategy. This formulation is particularly suitable when limiting peak actuator usage is critical. Similar to the $\ell_{1}$-norm, $\ell_{\infty}$-based control allocation problems can be cast as linear programs [Petersen_2005, Bodson_2011, Huang_2019].
It is worth noting that, for low-dimensional allocation problems or operating regimes in which actuator constraints remain inactive, some of the listed norm choices may yield very similar or in certain cases identical solutions. Pronounced differences typically arise in strongly constrained or highly over-actuated systems, where the geometry of the feasible control set interacts nontrivially with the selected norm.
Figure 4: Results of a comparative analysis illustrating the influence of different norm choices on control allocation with inactive and active input constraints.
To further illustrate the influence of the chosen norm on the solution of constrained control allocation problems, a simplified static example is considered and visualized in Figure 4. The problem involves a system with a two–dimensional control input vector $\mathbf{u}=[u_{1},\,u_{2}]^{T}$, whose admissible set is bounded by magnitude constraints $\mathbf{u}_{\min}\leq\mathbf{u}\leq\mathbf{u}_{\max}$ and indicated by the dashed red box in the figure. The control effectiveness matrix is defined as $\mathbf{B}=[1,\,-1]$, and the demanded virtual control input is set to $\nu=1$. The set of all control inputs satisfying the virtual control demand $\nu=\mathbf{B}\mathbf{u}$ forms a straight line in the $(u_{1},u_{2})$-plane and is shown as the dashed black line in Figure 4. Two representative cases are examined. In Case 1, the magnitude bounds are chosen as $\mathbf{u}_{\min}=[-2,\,-2]^{T}$ and $\mathbf{u}_{\max}=[2,\,2]^{T}$, such that the admissible region fully contains the unconstrained minima of the respective cost functions for the commanded $\nu$ of the case. Consequently, the input constraints are inactive, and the solution is determined solely by the norm-induced objective. In Case 2, the admissible region is deliberately reduced by selecting $\mathbf{u}_{\min}=[0,\,-1.5]^{T}$ and $\mathbf{u}_{\max}=[0.5,\,1.5]^{T}$, such that only a limited segment of the solution line remains feasible. In this case, the input constraints are active and directly influence the resulting allocation. For each case, three objective functions of the form $J=\lVert\mathbf{u}\rVert_{p}$ are considered, where the choice of $p$ corresponds to the $\ell_{2}$-, $\ell_{1}$-, or $\ell_{\infty}$-norm. The corresponding optimization problems are formulated using the appropriate numerical frameworks for each norm, as discussed in the previous part of this section. In all cases, the equality constraint $\nu=\mathbf{B}\mathbf{u}$ as well as the actuator magnitude limits are enforced as hard constraints within the optimization, as presented in Eq. (10). The resulting cost contours are visualized in Figure 4, enabling a direct geometric interpretation of the norm-induced allocation behavior.
For the $\ell_{2}$-norm, the cost contours are circular, and the optimal solution corresponds to the feasible point on the solution line that minimizes the Euclidean distance to the origin. In Case 1, this yields the solution $\mathbf{u}=[0.5,\,-0.5]^{T}$, resulting in a balanced distribution of control effort between the two inputs. In Case 2, the unconstrained minimum lies on the boundary of the admissible set, and the QP-based implementation is able to identify this solution, again resulting in $\mathbf{u}=[0.5,\,-0.5]^{T}$. In contrast, the $\ell_{1}$-norm yields diamond-shaped cost contours, corresponding to squares that are rotated by $45^{\circ}$ with respect to the coordinate axes in the $(u_{1},u_{2})$-plane. As a consequence, the associated optimization problem exhibits a sparsity-promoting behavior, tending to concentrate the control effort on a single actuator whenever this is compatible with the constraints. In Case 1, the $\ell_{1}$-based allocation yields the solution $\mathbf{u}=[1,\,0]^{T}$, where the entire control effort is assigned to $u_{1}$ while $u_{2}$ remains inactive. This behavior highlights the tendency of $\ell_{1}$-based approaches to reach the magnitude saturation of the most effective actuators at an earlier stage, which can be undesirable in applications where actuator saturation is problematic or where a balanced usage of available actuators is preferred. Although this effect can be reduced through appropriate conditioning and weighting of the optimization problem, as discussed in [Schierman_2004], the underlying numerical tendency remains inherent to the $\ell_{1}$-norm formulation. In Case 2, the active input limits significantly restrict the admissible solution space, such that the unconstrained $\ell_{1}$-minimum becomes infeasible. Nevertheless, the LP-based allocation routine is able to identify the same feasible solution as obtained with the other norms, namely $\mathbf{u}=[0.5,\,-0.5]^{T}$. This illustrates that, as long as the solution space is nonempty, all considered optimization formulations are capable of identifying admissible solutions, even when the cost function alone does not yield a unique or well-conditioned minimum. The $\ell_{\infty}$-norm leads, similar to the $\ell_{1}$-norm, to square-shaped cost contours but aligned with the coordinate axes. For the considered example, the resulting solution is identical to the $\ell_{2}$-based solution in both scenarios, yielding $\mathbf{u}=[0.5,\,-0.5]^{T}$ in Case 1 and Case 2. This behavior can be attributed to the well conditioned structure of the considered control effectiveness matrix, where both control inputs contribute with comparable magnitude to the virtual control demand. Under such conditions, minimizing the peak control effort does not introduce a strong bias toward a particular actuator, and the $\ell_{\infty}$- and $\ell_{2}$-based solutions coincide.
A notable distinction between the considered norm formulations arises from their smoothness properties. While the $\ell_{2}$-norm defines a smooth and continuously differentiable objective function, both the $\ell_{1}$- and $\ell_{\infty}$-norms are non-smooth and exhibit sharp edges in their level sets [harkegaard2003]. As demonstrated by the simplified example, these geometric properties directly influence the resulting allocation behavior. In particular, the $\ell_{1}$-norm exhibits a pronounced tendency to drive highly effective actuators into saturation, even in cases where alternative feasible solutions with more balanced actuator usage exist. Although this sparsity-promoting behavior can be beneficial in certain applications, it is generally undesirable in flight control systems where early saturation of individual actuators may reduce robustness or limit available control authority. The $\ell_{\infty}$-norm, on the other hand, in some cases yields solutions similar to those obtained with the $\ell_{2}$-norm when the control effectiveness of each actuator is comparable, as observed in the presented example. However, due to its non-smooth cost structure similar to $\ell_{1}$, $\ell_{\infty}$-based formulations may still lead to non-smooth allocation behavior and abrupt changes in the optimal solution when the active set changes [Frost_2009, Frost_2010]. In dynamic settings, such effects can result in rapid switching of actuator commands, input chattering and frequent operation near rate limits, which is undesirable from both a performance and energy conservation perspective. While the inclusion of explicit rate constraints and adjustments on the numerical optimization structure can mitigate these effects to some extent, the underlying tendency toward non-smooth allocation remains inherent to the formulation.
Based on these considerations, an $\ell_{2}$-norm-based formulation is adopted in this work for the GHGV-2 application. The $\ell_{2}$-norm promotes a balanced distribution of control effort in accordance with actuator effectiveness and exhibits favorable smoothness properties in the presence of actuator constraints. Building on this formulation, a variety of control allocation algorithms that leverage the $\ell_{2}$-norm have been proposed and applied in the literature. These approaches differ in how the control allocation problem is structured and simplified, and in how actuator limitations are addressed within the allocation process. The following subsection provides a brief overview of commonly used $\ell_{2}$-based control allocation algorithms and discusses their respective advantages and limitations in the context of hypersonic glide vehicle applications.
2.2 Common $\ell_{2}$-based Control Allocation Algorithms
@@ 2.2.1 Pseudoinverse-Based Control Allocation
The majority of the control allocation problems are solved by pseudoinverse-based control allocation (PICA) using the Moore-Penrose pseudoinverse, which implicitly solves the following unconstrained least-square problem for the allocation of the incremental virtual control input vector $\bm{\nu}$: